# Expert Router: Orchestrating Efficient Language Model Inference through Prompt Classification

### v1

@@ Abstract
Large Language Models (LLMs) have experienced widespread adoption across scientific and industrial domains due to their versatility and utility for diverse tasks. Nevertheless, deploying and serving these models at scale with optimal throughput and latency remains a significant challenge, primarily because of the high computational and memory demands associated with LLMs. To tackle this limitation, we introduce Expert Router, a system designed to orchestrate multiple expert models efficiently, thereby enhancing scalability. Expert Router is a parallel inference system with a central routing gateway that distributes incoming requests using a clustering method. This approach effectively partitions incoming requests among available LLMs, maximizing overall throughput. Our extensive evaluations encompassed up to 1,000 concurrent users, providing comprehensive insights into the system’s behavior from user and infrastructure perspectives. The results demonstrate Expert Router’s effectiveness in handling high-load scenarios and achieving higher throughput rates, particularly under many concurrent users.
@@
Large Language Models, Performance, Throughput
@@ I Introduction
Foundation models, particularly large language models (LLMs) like GPT-4 [1], Claude 3 [2], and Llama 2 [3] offer versatile applications across diverse industrial and scientific domains [4], such as the automotive domain [5], software development [6], physics [7], and medicine [8]. LLMs and their underlying transformer architecture [9] have been used in solving many natural language processing problems including sentiment analysis [10] and classification. More recently generative tasks such as summarization, text generation, and conversational systems are a dominant field of application. Many of these tasks demand multiple LLM inferences, e. g., to execute complex prompt sequences derived from different prompt engineering approaches [11], to utilize data integrations via retrieval augmentation generation, or to orchestrate agent-based workflows for optimal results [12].
The need to efficiently conduct inference and serve these models is critical to enable these applications. However, the deployment of LLMs presents significant challenges. This includes the complex model architecture and addressing their high computational and memory demands, which often necessitates specialized hardware like GPUs. For larger models, e. g., models beyond 70 billion parameters, it is required to partition the model across multiple GPUs utilizing model parallelism (e. g, tensor, and pipeline parallelism) or other optimization methods (e. g., pruning and quantization).
Smaller models can easily fit into the memory of a single GPU and do not require expensive multi-node tensor parallelism. Thus, such small models have recently received increased attention from industry and research. For example, Chinchilla models [13] prioritize training on more tokens rather than increasing the number of parameters, thus optimizing compute to achieve smaller yet higher-performing models. Other small language models, e. g., Phi-2, utilize high-quality training data [14]. Furthermore, using a Mixture of Experts (MoE), i. e. models that partition parameters into experts, each with unique weights, avoid the need to utilize all weights during training and inference [15] (e. g., Mixtral [16]). Other approaches utilize multiple independent smaller expert models and couple them hierarchically (e. g., Tabi [17]) or using other routing and coordination approaches.
Gururangan et al. [18] introduced the Cluster-Branch-Train-Merge (c-BTM) technique, which employs unsupervised clustering using k-means to partition a given dataset into distinct domains as the basis for multiple expert models that are trained in parallel. During inference, only experts relevant to the topic of the incoming prompt are utilized. We extend Gururangan’s work to new modern LLM models, such as LLama 2, and focus on characterizing the inference performance. For this purpose, we investigate C-BTM’s ensemble of expert approaches and compare them to the inference of larger parameter models using throughput and latency.
In this work, we introduce Expert Router, an inference system for expert model ensembles. Expert Router is designed for an ensemble of expert systems. It integrates with state-of-the-art inference systems, particularly TensorRT-LLM and Triton. Our modular system design is capable of integrating various models and inference systems. We evaluate Expert Router using extensive experimentation involving up to 1000 concurrent users and assess the system performance from user and system perspectives.
We summarize our contributions in this work as follows:
  • We introduce a highly modular architecture called Expert Router that allows us to orchestrate multiple expert models efficiently.
  • We investigate several configurations of Expert Router, including quantized and non-quantized state-of-the-art LLMs. These configurations are benchmarked against a baseline model on a diverse set of metrics grouped into user and system perspectives.
  • We conduct an extensive evaluation of all architectures with up to 1,000 concurrent users, in total producing over one billion response tokens to prove Expert Routers’ effectiveness in handling high-load scenarios.
II Related Work
Efforts to improve LLM serving can be grouped into two main categories: algorithmic innovations and system optimization [19]. In this section, we survey the techniques developed within these categories as well as for benchmarking these systems.
II-A LLM Algorithmic Innovations
Recently, MoE models have gained much attention due to their high performance in various benchmarks [20, 21, 22, 23, 16]. MoE models can be classified into two categories: integrated experts and ensembles of experts. An integrated MoE model uses a router network in specific layers to incorporate the output of distinct parameter groups, the experts. Ensembles of experts rely on separate models that are independently trained. All or a subset of these models are activated during inference, and the outcome—either the final output or the output activity of specific layers—is averaged.
An example of an integrated MoE model is Mixtral, which consists of eight seven-billion-parameter models [21]. It outperforms state-of-the-art classical LLM architecture in several benchmarks, such as LLama2, GPT-3.5, or Gemini Pro [21].
Gurungan et al. [18] propose an ensemble of expert approach called c-BTM. This technique involves training sparse language models on text corpora with minimal inter-node communication. It achieves this by clustering text corpora into related topics and training a dedicated expert language model for each cluster. Once trained, these models are collectively utilized for prediction tasks. We adopt the k-means clustering algorithm used and refined by Gururangan et al. to schedule incoming prompts in our architecture, leveraging their results for our model orchestration framework.
Another example of an ensemble of experts is OrchestraLLM [24]. This approach uses small and large language models for dialogue state tracking. OrchestraLLM’s dynamic routing mechanism for dialogue state tracking involves evaluating each input instance to decide the best-fit language model for processing. This process starts by calculating the semantic embedding of the input. Then, it compares this embedding with semantic representations from various expert pools using a cosine distance measure. The system then routes the input to the appropriate expert model based on a majority vote from these comparisons.
II-B LLM System Optimization
In addition to optimizing the core architecture of LLMs, improving the efficiency of inference calculations by optimizing the model deployment is possible. A key strategy in this domain involves modifying the memory storage format of model weights, activation, and user context, which is known as quantization [19]. Thereby, the general architecture is left unchanged, and only the data format is altered [25]. Various methods exist that allow transforming parameters that have been initially stored in FP32 or FP16 format to lower resolution storage formats such as FP8, INT8, INT4 or even INT2 [26, 27, 28, 29, 30]. For many of these Post Training Quantization (PTQ) techniques, it is possible to develop efficient GPU kernels [27]. This translates into reduced memory requirements and can lead to measurable latency reductions in certain settings [31].
Besides quantization techniques, the parallelization of models significantly influences the inference efficiency [32, 33]. A common approach called tensor parallelization distributes one LLM across multiple GPUs on the same node. Weight matrices are sliced and distributed across different GPUs depending on the implementation. Each GPU evaluates the corresponding matrix multiplication independently, and the final result can be derived with only little communication overhead. Another degree of parallelization is called pipeline parallelization, where different layers of the network are evaluated on different GPUs [34].
Wang et al. [17] proposed the ”Tabi” multi-level inference system for coordinating multiple small models and more complex LLMs using dynamic routing techniques. The core idea is to use less resource-intensive models for more straightforward queries instead of employing an LLM for every task.
The last layer in the system stack is formed by the software framework used to deploy the models. These programs often come with pre-set batching and scheduling techniques and optimized kernels for transformer architectures [19]. Examples are vLLM, ONNX Runtime, NVIDIA Triton, TorchServe, DeepSpeed Inference and HuggingFace text generation [35, 36, 37, 38, 39, 40].
One approach to optimize LLM serving specifically targeting the batching mechanism has been introduced by Agrawal et al. [41]. Their work segments the inference process into two distinct phases: prefill and decode. The prefill phase is optimized for parallel processing, fully engaging GPU computational power, but at the expense of higher latency. Conversely, the decode phase processes output tokens sequentially, not fully utilizing the available GPU compute. This creates the challenge of balancing high throughput and low latency in LLM serving systems. Agrawal et al. have developed Sarathi-Serve, an inference scheduler that employs chunked-prefills and stall-free scheduling to address this issue.
@@ II-B LLM System Optimization
II-C Benchmarking of Inference Architectures
Performance characteristics can be assessed from two perspectives: system and user. On the system side, an important metric is throughput, defined as the number of tokens generated by the LLM per second [19]. From the user’s viewpoint, several metrics beyond throughput are relevant, including first-token response time, inter-token latency, and total processing time [42].
As stated by Miao et al., creating a holistic evaluation framework that measures the performance of LLM inference systems is complicated due to the large configuration space of the underlying architectures [19]. Recently, MLPerf has added a GPT-J test scenario to their benchmarking suite where it is possible to evaluate the number of samples and queries per second [43, 44]. Another benchmarking framework that has lately received significant attention from the community is LLMPerf, which Anyscale developed. Their platform allows measuring key characteristics such as first token response time, throughput, inter-token latency etc. [45, 42].
Our work builds upon these approaches for model optimization and deployment, directly addressing the challenge of coordinating multiple expert models efficiently. For this, we are employing the k-means cluster algorithm developed by Gurungan et al. for routing user prompts to the matching expert model [18]. Therefore, our approach differs from the traditional MoE methodology by not averaging any activation outputs but rather activating models individually. Each model in our approach is fully autonomous, generating its own distinct output. This architectural choice is motivated by the flexibility it introduces, allowing for the extension to appropriate domains. Another key strength is its ability to incorporate many optimization techniques, such as system optimizations and algorithmic innovations. Its modular design ensures that components can be seamlessly integrated or updated. Our work differentiates itself from a classical data parallel architecture, as we do not statically but dynamically route incoming prompts based on their topic. To thoroughly evaluate our architecture, we employ many of the aforementioned metrics.
@@ III System Architecture
This section presents Expert Router’s architecture, which enables efficient and scalable LLM inference. The architecture is highly modular, i. e., individual components can be replaced or upgraded, allowing for continuous and systematic improvements. For example, the architecture enables LLMs to be individually updated without necessitating a complete system update. Expert Router can be deployed in a heterogeneous computing environment irrespective of the hardware components’ manufacturing era.
Figure 1 illustrates the system architecture. The routing gateway is the central component of the system. It manages request distribution to the inference servers. Specifically, we utilize NVIDIA Triton [37] and TensorRT-LLM [46] to execute the inference tasks. The system can easily be scaled on every layer, e.g., utilizing a load balancer like NGNIX before the routing gateway or adding new Triton server instances. It utilizes the gRPC protocol [47] for communication between these components. Expert Router can achieve high throughputs through this design while providing a flexible solution for LLM inference.
@@ III-A Routing Gateway
The routing gateway acts as an intermediary; it receives and distributes requests across inference servers. This functionality is also referred to as reverse proxy. After receiving a new inference request, the routing gateway extracts the user prompt. It sends it to the query classification (Step 3/4 in Figure 1), where a k-means clustering algorithm is used to determine the target cluster of the prompt. This mechanism is directly adopted from the system developed by Gururangan et al. [18].
We follow the original paper’s preprocessing steps: excluding stop-words and substituting numerical tokens. The clustering process transforms the prompt text data using a TF-IDF vectorizer from the scikit-learn library [48]. The TF-IDF vectors are then condensed to 100 dimensions via a singular value decomposition. This dimensionality reduction is followed by a normalization step by subtracting the mean and scaling the vectors to unit variance [18]. We employ the same custom PyTorch implementation, including the weights of the k-means model published by [18], which has been trained on a selected portion of the C4 dataset [49].
Expert Router uses the pre-trained k-means model to classify the prompt into an appropriate cluster when receiving a new prompt. The routing gateway then directs the prompt to the designated language model hosted on a Triton server instance. To increase prompt classification throughput, we are deploying 16 routing gateway instances in combination with an NGINX load balancer that distributes the incoming requests in a round robin mechanism [50] (Step 1/2 and 7/8 in Figure 1).
@@ III-B Inference Backend and Triton Server
Following an overview of large language model (LLM) inference techniques, we discuss the particular inference backend employed by Expert Router: the GPU-accelerated NVIDIA Triton server.
@@
We rely on advanced inference techniques, particularly optimization for the self-attention mechanisms and batching. Self-attention enables every token of the input sequence to be evaluated against all preceding tokens, thereby creating a contextual representation. This process starts by converting input tokens into Query (Q), Key (K), and Value (V) vectors through trained weights. Next, it calculates the dot products between all Q and K vectors and applies a softmax function. The outcome of this calculation is combined with the V vectors through another dot product to produce a weighted sum. This weighted sum models the context between tokens.
Multi-head attention enables transformers to leverage multiple attention mechanisms simultaneously, with each ”head” focusing on distinct representation subspaces. These heads independently compute attention weights for the Q, K, and V vectors. By aggregating the outputs of these heads, multi-head attention allows the capture of a broader range of relationships within the data. Further, this allows for parallelized examination of input sequences [9].
A critical optimization within this process is the use of a KV-cache. This cache stores the K and V vectors from previous computations, eliminating the need for redundant recalculation [41].
In inference scenarios of LLMs, there are two main processing stages. The first stage is known as the prefill phase. During this phase, the system processes the entire input sequence to produce the first output token, simultaneously populating the KV-cache with K and V vectors derived from the input. This stage primarily relies on matrix-matrix parallelization, which can be efficiently parallelized [41].
Following the prefill phase, new tokens are generated autoregressively, leveraging the precomputed activations stored in the KV-cache. This is called the decode phase. The decoding phase is computationally less demanding, often underutilizing the GPU [41]. To address this inefficiency, employing batching techniques becomes essential for optimizing LLM service, especially in high-load scenarios.
@@
We use NVIDIA’s Triton server as an inference backend. The Triton server starts with the inference process after receiving the prompt from the routing gateway (Step 5 in Figure 1). Each model is hosted on a dedicated Triton server, with each server assigned to its own GPU. This configuration allows the system to individually address models via distinct ports, each linked to a specific cluster identified by the k-means clustering algorithm. This architectural choice eliminates the need for communication between different models or GPUs, as each Triton server is equipped with its scheduling and in-flight batching mechanisms. We use each Triton server in streaming mode, i. e., tokens generated by the model are instantly transmitted back to the routing gateway, avoiding any buffering of tokens on the system side (Step 6 in Figure 1). The routing gateway then forwards these tokens to the end-users through the NGINX load balancer (Steps 7/8 in Figure 1).
To enhance inference performance, we utilize TensorRT-LLM, a specialized framework designed to build and deploy LLM engines. TensorRT-LLM is optimized for NVIDIA hardware, ensuring that state-of-the-art models such as LLama2, Mixtral, and Gemini can operate efficiently.

### v2

# Performance Characterization of Expert Router for Scalable LLM Inference

@@ Abstract
Large Language Models (LLMs) have experienced widespread adoption across scientific and industrial domains due to their versatility and utility for diverse tasks. Nevertheless, deploying and serving these models at scale with optimal throughput and latency remains a significant challenge, primarily because of LLMs’ high computational and memory demands. Specialized models optimized for specific tasks can be combined through a routing mechanism to address these challenges, creating a modular inference system. This paper introduces Expert Router, a scalable routing architecture that directs prompts to specialized expert models. We characterize multiple Expert Router configurations, including different LLama 3 models with quantized and non-quantized weights under up to 1,000 concurrent users. Our findings reveal that Expert Router introduces minimal latency overhead, with the configuration of expert models being a dominating factor in performance outcomes. High-parameter expert models deliver stable throughput and latency under moderate concurrency levels. In contrast, smaller expert models maintain competitive performance across a wider range of concurrent users compared to tensor-parallelized baseline models. This highlights the potential of Expert Router for efficient and scalable LLM deployment.
@@
Large Language Models, Inference, Performance, Routing
@@ I Introduction
Foundation models, particularly large language models (LLMs) like GPT-4 [1], Claude 3 [2], and Llama 3 [3] offer versatile applications across diverse industrial and scientific domains [4], such as the automotive domain [5], software development [6], physics [7], and medicine [8]. These models are based on the transformer architecture [9], which uses the self-attention mechanism to calculate relationships within the input sequence and can solve many natural language processing problems, including sentiment analysis [10] and classification [11]. More recently, generative tasks such as text summarization, text generation, and conversational systems have become important. Many of these tasks demand multiple LLM inferences, e. g., to execute complex prompt sequences derived from different prompt engineering approaches [12], to utilize data integrations via retrieval augmentation generation, or to orchestrate agent-based workflows for optimal results [13].
The need to efficiently conduct inference and serve these models is critical to enable these applications. However, the deployment of LLMs presents significant challenges. This includes the complex model architecture and addressing their high computational and memory demands, often necessitating specialized hardware like GPUs. For larger models, e. g., models beyond 70 billion parameters, it is required to partition the model across multiple GPUs utilizing model parallelism (e. g. tensor and pipeline parallelism) or other optimization methods (e. g. pruning and quantization).
A different approach to larger models is the utilization of expert models that are tailored to perform well in specific tasks. Combining multiple specialized models through a routing mechanism, which directs queries to the most suitable LLM and creates a single inference system that improves overall answer quality. Recent studies have shown that routing-based inference systems, when configured with smaller LLMs, can outperform larger models regarding response quality [14].
However, beyond improvements in response quality, routing inference systems also impact critical performance characteristics such as latency and throughput. Thereby, the routing architecture should maintain stable latency and throughput, even when operating under the load of multiple concurrent users. Furthermore, these systems should be capable of working with different model configurations, enabling them to utilize specific model strengths to meet different performance requirements. Additionally, routing systems should be scalable and adaptable, functioning efficiently in different computing environments.
In this work, we introduce Expert Router, a scalable routing inference architecture that combines multiple expert models into one system. Expert Router utilizes a technique called c-BTM introduced by Gururangan et al. [15]. It employs unsupervised clustering using k-means to partition a given dataset into distinct domains as the basis for multiple expert models trained in parallel. We adopt this architecture to create an embarrassingly parallel inference pipeline where incoming prompts are routed to the correct expert model. To evaluate Expert Router, we conduct extensive experiments with up to 1,000 concurrent users, recording the timestamp of each output token to generate the relevant system and user-centric metrics. This data provides insights into the characteristics of Expert Router, allowing us to assess the impact of routing on latency and throughput. We summarize our contributions in this work as follows:
  • We present Expert Router, a flexible and scalable architecture that efficiently combines multiple expert models, enabling dynamic routing and seamless integration of various model configurations to optimize LLM inference.
  • We evaluate multiple configurations of Expert Router, including quantized and non-quantized LLMs, benchmarking them against baseline models across user and system metrics.
  • Based on our findings, we provide recommendations for optimizing routing-based inference systems, focusing on balancing latency, throughput, and resource allocation.
II Related Work
Multiple studies [16, 17, 18, 19, 20] have explored LLM routing strategies to optimize inference costs. These studies route simple requests to weaker LLMs and direct more complex tasks to more capable models, such as GPT-4. For instance, Wang et al. [21] proposed an inference system that uses dynamic routing techniques to distribute queries among multiple small models and complex LLMs. This approach uses less resource-intensive models for simpler queries rather than relying on an LLM for every task.
Other studies [22, 23, 24] have investigated training a routing function to direct prompts to the most suitable LLM based on their capabilities. As an example, Shnitzer et al. [14] are using scoring or ranking models to evaluate the strengths and weaknesses of different LLM candidates. The study identifies that it is possible to outperform a larger model by routing to smaller LLMs.
In the field of LLM performance evaluation, recent studies have established metrics that allow benchmarking and comparing different deployment methods. Most notably, Reddie et al. [25] and MLPerf have released a benchmark framework and reference implementations to evaluate inference backends based on a LLama 2 70B model [26]. The group defines and describes metrics such as the Time to First Token (TTFT), time per output token (TPOT) and throughput in token per second (token/s). These metrics are also used by other publications to compare the performance of hardware and deployment architectures [27, 28, 29, 30]. Additionally to the above streaming related quantities, multiple studies [31, 32, 33] use classical serving metrics such as the total latency or the p99 latency.
LLM architectures can undergo different runtime conditions depending on factors such as the available hardware, number of requests and deployment hyperparameters. To test these different settings it is necessary to measure the change of the mentioned metrics under varying load. Previous work has investigated how factors such as the number of input tokens, batch size and number of concurrent users can affect these metrics [33, 34, 35]. In the aforementioned MLPerf benchmark, the simulated queries cover a broad spectrum of input token counts, which reflects varying load scenarios [26]. Agrawal et al. introduced Viduar: A Large-Scale Simulation Framweork for LLM Inference [30]. In their framework, different workloads can be selected which are based on different input datasets, request types and arrival rates.
Our work explores the effects of model routing on user and system metrics, particularly under a high number of concurrent users. While previous research has shown that routing can improve answer quality, our study focuses on how it influences different latency metrics and throughput values.
@@ III System Architecture
This section presents Expert Router’s architecture, which enables efficient and scalable LLM inference. The architecture is highly modular, i. e., individual components can be replaced or upgraded, allowing for continuous and systematic improvements. For example, the architecture enables LLMs to be individually replaced without necessitating a complete system update. Expert Router can be deployed in a heterogeneous computing environment irrespective of the hardware components’ manufacturing era.
Figure 1 illustrates the system architecture. The routing gateway is the central component of the system. It manages request distribution to the inference backend. Specifically, we utilize NVIDIA Triton [36] and TensorRT-LLM [37] to execute the inference tasks. The system can easily be scaled on every layer, e.g., utilizing a load balancer like NGNIX before the routing gateway or adding new Triton server instances. It utilizes the gRPC protocol [38] for communication between these components.
@@ III-A Routing Gateway
The routing gateway acts as an intermediary; it receives and distributes requests across inference servers. This functionality is also referred to as reverse proxy. After receiving a new inference request, the routing gateway extracts the user prompt. It sends it to the query classification (Step 3/4 in Figure 1), where a k-means clustering algorithm is used to determine the target cluster of the prompt. This mechanism is adopted from the system developed by Gururangan et al. [15].
We follow the original paper’s preprocessing steps: excluding stop-words and substituting numerical tokens. The clustering process transforms the prompt text data using a TF-IDF vectorizer from the scikit-learn library [39]. The TF-IDF vectors are then condensed to 100 dimensions via a singular value decomposition. This dimensionality reduction is followed by a normalization step by subtracting the mean and scaling the vectors to unit variance [15]. We employ the custom PyTorch implementation, including the weights of the k-means model published by [15], which has been trained on a selected portion of the C4 dataset [40]. Working with the reduced-dimensional input sequence is especially efficient in inference scenarious, as the classification process only scales with the number of concurrent users but not with the length of the input requests.
Expert Router uses the pre-trained k-means model to classify the prompt into an appropriate cluster when receiving a new prompt. The routing gateway then directs the prompt to the designated language model hosted on an inference backend instance. Unlike c-BTM, which dynamically selects and combines multiple experts during inference, our approach streamlines the process by routing each request directly to a single, most relevant model, reducing computational overhead during real-time inference. To increase prompt classification throughput, we are deploying 16 routing gateway instances in combination with an NGINX load balancer that distributes the incoming requests in a round robin mechanism [41] (Step 1/2 and 7/8 in Figure 1).
@@ III-B Inference Backend and Triton Server
Following an introduction into LLM inference techniques, we discuss the particular inference backend employed by Expert Router: the GPU-accelerated NVIDIA Triton server.
@@
We rely on advanced inference techniques, particularly optimized for the self-attention mechanisms and batching. Self-attention enables every token of the input sequence to be evaluated against all preceding tokens, thereby creating a contextual representation. This process starts by converting input tokens into Query (Q), Key (K), and Value (V) vectors through trained weights. Next, it calculates the dot products between all Q and K vectors and applies a softmax function. The outcome of this calculation is combined with the V vectors through another dot product to produce a weighted sum. This weighted sum models the context between tokens. A critical optimization within this process is the use of a KV-cache. This cache stores the K and V vectors from previous computations, eliminating the need for redundant recalculation [27].
In inference scenarios of LLMs, there are two main processing stages. The first stage is known as the prefill phase. During this phase, the system processes the entire input sequence to produce the first output token, simultaneously populating the KV-cache with K and V vectors derived from the input. This stage primarily relies on matrix-matrix computations, which can be efficiently parallelized [27].
Following the prefill phase, new tokens are generated autoregressively, leveraging the precomputed activations stored in the KV-cache. This is called the decode phase. The decode phase is computationally less demanding, often underutilizing the GPU [27]. To address this inefficiency, employing batching techniques becomes essential for optimizing LLM service, especially in high-load scenarios.
@@
Expert Router utilizes multiple independent models that are each processing their own batch of input requests. These multiple batches introduce another level of asynchrony in the inference system leading to distinct latency and throughput behaviours.
@@
We use NVIDIA’s Triton server as an inference backend. The Triton server starts with the inference process after receiving the prompt from the routing gateway (Step 5 in Figure 1). Each model is hosted on a dedicated Triton server, with each server assigned to its own GPU. This configuration allows the system to individually address models via distinct ports, each linked to a specific cluster identified by the k-means clustering algorithm. This architectural choice eliminates the need for communication between different models or GPUs, as each Triton server is equipped with its scheduling and in-flight batching mechanisms. We use each Triton server in streaming mode, i. e., tokens generated by the model are instantly transmitted back to the routing gateway, avoiding any buffering of tokens on the system side (Step 6 in Figure 1). The routing gateway then forwards these tokens to the end-users through the NGINX load balancer (Steps 7/8 in Figure 1). To enhance inference performance, we utilize TensorRT-LLM, a specialized framework designed to build and deploy LLM engines.